OpenStudy (anonymous):

Given the exponential equation 3x = 243, what is the logarithmic form of the equation in base 10?

OpenStudy (anonymous):

I got B as my answer

OpenStudy (anonymous):

Am i correct

OpenStudy (solomonzelman):

\(\large\color{black}{ 3^x=243}\) gives \(\large\color{black}{ \log 243=x \log }3\) \(\large\color{black}{ \log 243=x \log }3\) \(\large\color{black}{ \log 243 \div \log3=x }\) or \(\large\color{black}{ \log _3243=x }\)

OpenStudy (solomonzelman):

Yes you are

OpenStudy (anonymous):

Thanks once again!

OpenStudy (solomonzelman):

I would just say that 243 = 3^5 if I was to solve it. No trig solutions to algebra

OpenStudy (solomonzelman):

yw

OpenStudy (adilalvi):

If we take the log (base 10) of both sides we have: log (base 10) 3^x = log (base 10) 243 Using the law that says log a^b = b*log a we have: x*log (base 10) 3 = log (base 10) 243 x = [log (base 10) 243]/[log (base 10) 3] x = [log (base 10 (2.43*100)]/[log (base 10) 3] Using the law that says log (a*b) = log a + log b: x = [log (base 10) 2.43 + log (base 10) 100]/[log (base 10) 3] Since the log (base 10) of 100 = 2 (remember 10^2 = 100 so the log (base 10) of 100 = 2): x = [(log (base 10) 2.43) + 2]/[log (base 10) 3. Look up the base 10 logs for 2.43 and 3 and finish the calculation.

OpenStudy (solomonzelman):

wow... if I didn't know how to do this on my own, this would scare me so much !

OpenStudy (adilalvi):

what you mean this is wrong?